THE HOT SUMMER OF HITRAN2008

\underline{I. E. GORDON, L. S. ROTHMAN}, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA.

The HITRAN2008^{a} database is expected to be officially released at the end of this summer. The ”β-version” of the new compilation will be assembled before the HITRAN meeting which immediately follows Ohio symposium. At the HITRAN meeting, the results of the quality assessments of the ”β-version”, as well as suggested ”last-minute” additions, will be discussed. The new data will be comprised of the line lists that are already available on the web site (http://www.cfa.harvard.edu/hitran/updates.html) and some of the forthcoming updates. The line-by-line lists for almost all of the 39 HITRAN molecules will be updated in comparison with the previous compilation (HI-TRAN2004^{b}). The caliber of the updates ranges from corrections to the representations of quantum assignments to complete replacements of the lists. In addition, 2 new molecules, BrO and CH_{3}Br, will be included. Some of the most important updates, such as for water vapor, CO_{2}, O_{3}, the oxygen A-band, etc, will be presented.

aThe HITRAN database is supported by the NASA Earth Observing System (EOS) under the grant NAG5-13524.

bL. S. Rothman et al., JQSRT 96, 139 (2005).